# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v1

@@ 5.1 Assuming a partial Riemann hypothesis
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
@@
Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound $|\zeta(s)|\leq\zeta(\sigma)\leq\sigma/(\sigma-1)$, which is decreasing for all $\sigma>1$. Together with the trigonometric inequality from (21), we see that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$
@@ 5.3 Combining both cases $t\leq H$ and $t\geq H$
Second, a computation from the proof of [2, Proposition A.2], implemented in interval arithmetic, states
@@
which is an improvement over what one gets from [4, Theorem 1.1].
@@ 6 Concluding Remarks
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=W_{0}$, the condition (34) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (22), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [24, §14], [21], [22], and [5] for more details).

### v2

@@ 5.1 Assuming a partial Riemann hypothesis
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$ (47)
@@
Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound
$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},\qquad(\sigma>1),$ (48)
where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Together with the trigonometric inequality from (23), we see that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$ (49)
@@ 5.3 Combining both cases $t\leq H$ and $t\geq H$
Second, a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states
@@
which is an improvement over what one gets from [5, Theorem 1.1].
@@
5.4 An asymptotic improvement
In this section we prove an asymptotic improvement over the bounds given earlier in Section 5, at the expense of a worse constant. For the sake of clarity, we first outline the general idea here, before proving the next two theorems.
Assume we have the bound
$\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
for $t\geq t_{0}$ and $1-\delta\leq\sigma\leq 1+\delta_{1}$, where $\delta=d/\log t$ and $\delta_{1}=d_{1}/\log t$, as before. Then from (45), we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
$\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((d_{1}+d)C\right).$ (53)
Notice that since the exponential factor is $O(1)$, the overall order is determined by the first factor.
With the same argument involving the trigonometric inequality (22),
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4},$ (54)
and at the expense of a worse constant, we can estimate the second factor in (54) differently from the previous sections. We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$. Along with Lemma 9, this leads to
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq(C_{4}\log t)^{3/4}(C_{3}(\log t)^{2/3})^{1/4},$
which applied to (5.4) ultimately furnishes the bound
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq C_{5}(\log t)^{11/12}.$
Theorem 6.
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log t_{0}\leq 1$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{RH(T)}(d,d_{1},t_{0})(\log t)^{11/12},$
where
$Y^{\prime}_{RH(T)}$ $(d,d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{4}$ defined in Theorem 4, $B_{t_{0}}$ defined in Theorem 2, and $C_{3}$ defined in Lemma 6.
Proof.
The proof is identical to that of Theorem 4, except that for $\sigma\leq 1+\delta_{1,t}$, by Lemmas 6 and 9, we can replace (47) with
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|$ $\leq|\zeta(1+\delta_{1,t})|^{3/4}|\zeta(1+\delta_{1,t}+2it)|^{1/4}$
$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}.$
Meanwhile, for $\sigma\geq 1+\delta_{1,t}$, the same argument with (22) and (48) leads to
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
in place of (49). ∎
Theorem 7.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{H}(d,d_{1})(\log t)^{11/12},$
where
$Y^{\prime}_{H}$ $(d,d_{1})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log H}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log H}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{5}$ defined in Theorem 5, and $C_{3}$ defined in Lemma 6.
Proof.
The proof is similar to that of Theorem 6, except we do not assume $RH(T)$. Then the process follows the proof of Theorem 5, making similar changes as with the earlier case $t\leq H$ in Theorem 6. ∎
Combining Theorems 6 and 7, we prove Corollary 5, a result which holds for all $t\geq t_{0}\geq 5$.
Proof of Corollary 5.
The proof is analogous to that of Corollary 4, where the final result is obtained by optimising over the appropriate parameters.
For the result on $\sigma=1$, the limiting region is again $2\exp(e^{2})\leq t\leq H$, where we have
$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12},$
with $d_{1}=0.72042$. ∎
$t_{0}$ $Y^{\prime}(W_{0},t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
$10$ $21.123$ $0.307587$ $W_{0}$ $47.054$ $0.170052$
$10^{2}$ $11.643$ $0.542829$ $6$ $43.568$ $0.173052$
$10^{3}$ $9.842$ $0.648796$ $7$ $38.179$ $0.178401$
$10^{4}$ $8.885$ $0.725024$ $8$ $34.753$ $0.182368$
$10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
$10^{7}$ $7.425$ $0.885013$ $10$ $30.668$ $0.187861$
$10^{9}$ $6.907$ $0.960526$ $11$ $29.356$ $0.18984$
$10^{11}$ $6.545$ $1.021396$ $12$ $28.325$ $0.191481$
$H$ $6.340$ $1.059541$ $13$ $27.494$ $0.192868$
Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places). Each entry is valid for $t\geq t_{0}$.
@@ 6 Concluding Remarks
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=W_{0}$, the condition (36) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (24), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [25, §14], [23], [22], and [6] for more details).
Even in the unconditional results used, there is room for improvement. As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large. In the case of $\sigma=1$, this happens for $t\geq\exp(161975)$. This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 5.
Finally, we conclude with a discussion on the trigonometric inequality (22) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$. Under certain conditions, this method is superior to using the trivial bound. In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets. Surprisingly, the same cannot be said in our setting (with the exception of results in Section 5.4), and the classical polynomial we used is actually optimal. Take for instance Lemma 12, which is used in Theorems 4 and 5. Here the overall order of the bound is $O(\log t)$, and in this scenario the article [18] has an in-depth discussion on why the classical polynomial is best.
On the other hand, in Theorems 6 and 7, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse. This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.